Professor Sprout decides to leave Hogwarts and run a business of mandrakes which have commercial due to its health restorative properties. She reads a research article in the Daily Prophet that provides the following production function (as a function of labor): ?(?) = ??? + ??? ? − ?. ?? ? Before she enters into business, she asks you to help her to figure out if it is feasible for her to enter in the market with 14 workers. I. (1 pt.) Find the average product of labor [APL] and marginal product of labor [MPL] formulas. II. (2 pts) Graph the total product curve, indicate the inflection point and the possible maximum value of production III. (2 pts) Graph the MP and AP curves and verify that AP = MP when AP is at its maximum IV. (1 pt.) What would be your advice? Would you allow her to enter into business or not? Why or why not?
Answer #1

We use the standard formulae of APL and MPL to calculate them using the given production function.

We advice Professor Sprout to undertake the business as it is having positive production and APL isn't maximum at 14 workers yet.

- Glven: We have the production function: 8 - 10L + 2042 -0.6 I we know that APL = 8 and MPL = da production Hence, substitut

11 Total Product Curve is . 8 = 10L + 2012 -0.623 8 - LC 10 +20L -0.614) If we try to factorise this further, we get DEL (L -AK Inflection point (33.8261,0 ) At this inflection point, we And that do = 0 as slope of curve is o. 10 + 40L - 1.8120 .. LSince we know that Loo at the Inflection point, we can say that. L = 22.4694 is the inflection point 0 (L=22.4694) = 10 (22.4MPL (L- 503) = 10 + ftoxs) -8% MPL = 176.6667 APL CL-50%) 10+ Boxse) – 6.6 xsongo APL = 176.6667 -> MPL - APL @ Alls maximumAt L = 14, 8 = 10 (14) +20(14)2 -0.644) 8= 2413.6 units Also, of is maximum @ 22.4694 workers Hence, although it is feasible